Abiodun and taiwo share 735 naira between them so that abiodun gets 95 naira more than taiwo.Find how much money each gets
Answer by stanbon(75887)

Abiodun and taiwo share 735 naira between them so that abiodun gets 95 naira more than taiwo.Find how much money each gets
-----------------
Equations:
A + T = 735 naira
A = T + 95
-------
A + T = 735
A - T = 95
--------------
Add and solve for A::
2A = 830
A = 415 naira (amt Abiodun has)
T = 735-415 = 320 naira (amt Taiwo has)
----------
Cheers,
Stan H.
---------